OpenStudy (anonymous):

Intro to Imaginary and Complex Numbers: 21 - (sqrt.) -16 -24 + (sqrt.) -81 16 + (sqrt.) -169 could anyone tell me the answers and show the work?

OpenStudy (anonymous):

B ur lucky im givong u anserws

OpenStudy (anonymous):

out of my entire worksheet, these three are the ones I'm stuck on

OpenStudy (anonymous):

yea they are pretty hard

OpenStudy (anonymous):

not really I'm just stuck on these three because it's hard to simplify two different set of numbers with different factors, guess I'm just unsure how to write it out

OpenStudy (anonymous):

and what the proper work would be to show how I got the answer

OpenStudy (anonymous):

yea wll good luck and btw ur super pretty

OpenStudy (anonymous):

can u give me a medal

OpenStudy (anonymous):

\[-24+\sqrt{-81}=-24+\sqrt{81\times -1}=-24+9\sqrt{-1}=-24+9\sqrt{\iota^2}=-24+9\iota \]

OpenStudy (anonymous):

@ kamman1260 you didn't halp me just replied on my post so I don't know why I would (not to be mean) @ surjithayer the last bit of the equation didn't show up I don't know

OpenStudy (anonymous):

\[-24+9\sqrt{\iota^2}=-24+9\iota \]

OpenStudy (anonymous):

\[write \sqrt{-a}=\sqrt{a \times -1}\]

OpenStudy (anonymous):

so what about the other two?

terenzreignz (terenzreignz):

See, in a nutshell, when you're taking the square root of a negative number, just get the square root of the absolute value and then, affix i. \[\Large \sqrt{-a^2}=|a|i\] \[\Large \sqrt{-100}= 10i\] \[\Large \sqrt{-25}=5i\] I'm sure you get the idea :)

OpenStudy (anonymous):

I understand that, I have a whole worksheet, I'm just stuck on those three. I figured out number 9, but number 7 and 8 are still wrong. .. along with a whole bunch I just figured out from my results on the quiz.

terenzreignz (terenzreignz):

Just so you know, I have no idea which is which :) Tell me which you got wrong and what your wrong answers were... let's troubleshoot ^_^

OpenStudy (anonymous):

these are the answers I got wrong, give me a minute to post the questions for them

OpenStudy (anonymous):

5) 3 (sqrt.) -9 6) 2 (sqrt.) -16 7) 21 - (sqrt.) -16 8) -24 + (sqrt.) -81

terenzreignz (terenzreignz):

Just so you know, don't make a habit of using attachments that have to be downloaded... try to make your content as accessible as possible... Now let's see...

terenzreignz (terenzreignz):

Too small. Just type in the questions that correspond to the answers, lol

OpenStudy (anonymous):

14) 12i (to the power of) 25 15) 8i (to the power of) 15 16) 24i (to the power of ) 20 17) i (to the power of) 255

OpenStudy (anonymous):

Didn't know you had to download the content, on my computer I can view them before downloading. Just try zooming in.

terenzreignz (terenzreignz):

The point, Lilly :) Just type in ONE question which you got wrong, and your answer to that question.

OpenStudy (anonymous):

I figured out three of the 4 I posted, so I'll let you know how it goes until you see what you can do if you can help me.

OpenStudy (anonymous):

nevermind, i failed it anyway I'm out of attempts but thank you for the help! Math is just not my best subject is all

terenzreignz (terenzreignz):

That's too bad. I'd say better luck next time, but I'd rather you don't rely on luck :) Just do your best :)